The following data were obtained from a survey of college students. The variable X represents the number of non-assigned books read during the past six months.

x            0 1 2 3 4 5 6

P (X=x) 0.55 0.15 0.10 0.10 0.04 0.03 0.03

Find P(1 < X < 5). Place your answer, rounded to two decimal places in the blank. For example, 0.56 would be a legitimate entry.
Answer #1

P(1<X<5)=p(x=2)+p(x=3)+p(x=4)

= 0.10+0.10+0.04

= 0.24

hence answer is 0.24

(RE) О: 15 010 2. 3 Lt S O:16 03 IBM Business Analytics Lab at Lloyd Business School O-24